OpenStudy (anonymous):

PLZ help !!! can someone plz help me solve this question. To register for online banking services at a local credit union, you need to select a four-digit personal identification number, or pin code. The digits cannot be repeated. How many ways are there to arrange four digits out of the numerals 0 through 9 to create a code?

OpenStudy (anonymous):

[10 * 9 * 8 * 7] = 5040

OpenStudy (anonymous):

(x)(x-1)(x-3) or 10*9*8*7=5040

OpenStudy (anonymous):

is this how we do it ? nᴾr = n! / (n - r)! 10ᴾ4 = 10! / (10 - 4)! 10ᴾ4 = (10*9*8*7*6*5*4*3*2*1) / ( 6*5*4*3*2*1) 10ᴾ4 = 3628800 / 720 10ᴾ4 = 5040

OpenStudy (anonymous):

is 5040 the pin code ? but its say no digit can be repeated right ? but in the answer 0 i repeated

OpenStudy (anonymous):

5040

OpenStudy (anonymous):

lets say x is the chosen digit.....you have 10 choices for the first one, if you take that out because it is unusable, you have 9 choices for the second, 8 for the third, etc....so the total number of combinations would be 5040

OpenStudy (anonymous):

In the answer the digits can repeat

OpenStudy (anonymous):

Because it's the number that many time 10 9 8 7 can be arranged

OpenStudy (anonymous):

Many different times*

OpenStudy (anonymous):

E.g. Code can 98710 or 87910 .... 5040 different codes

OpenStudy (anonymous):

oh ok thank you so much i understand now. is the process i have written above the correct way to right it ?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

ok thank you